OpenStudy (anonymous):

which set represents the complete list of solutions of the equation log(10x)+log(x-3)=2? How do u do this?

OpenStudy (asnaseer):

do you know your log rules? e.g. log(a) + log(b) = log(ab)

OpenStudy (anonymous):

Yes

OpenStudy (asnaseer):

so use that rule to first simplify the left-hand-side of the equation.

OpenStudy (anonymous):

so you get 10x^2-30x=2. Do you subtract the 2 to the left side and then use the quadratic formula?

OpenStudy (asnaseer):

not quite, what you get first is: log(10x^2-30x)=2

OpenStudy (asnaseer):

is the log function here base 10 or is it the natural log?

OpenStudy (anonymous):

What do you mean by natural log?

OpenStudy (asnaseer):

natural log of x would be \(log_e(x)\) base 10 log of x would be \(log_{10}(x)\)

OpenStudy (asnaseer):

right, so next you need to use the definition of what a log is, i.e.:\[\text{if }\quad log_{10}(x)=y\]then\[x=10^y\]

OpenStudy (asnaseer):

can you see what to do next?

OpenStudy (anonymous):

10^(10x^2-30x)=2 ?

OpenStudy (asnaseer):

no - try again - look carefully at the definition of what a log represents

OpenStudy (anonymous):

Ohh sorry it's 100=10x^2-30x

OpenStudy (asnaseer):

perfect! - so now you just need to solve this quadratic equation and you are home and dry!

OpenStudy (anonymous):

Thank you for taking your time to help!:)

OpenStudy (asnaseer):

no problem - you're welcome